# Semilinear spaces of matrices with transitivity, rank or spectral constraints

### v1

@@ 5.2 The transitive rank
Denote by $r^{\prime}$ the rank of $\mathbf{C_{1}}$: it is also the transitive rank of $\mathcal{S}^{\prime}$. If $r^{\prime}<(n-1)d-1$, then by induction $\dim\mathcal{S}^{\prime}<\alpha(n-1,d)-(n-1)+1+\varepsilon$ and hence
$\dim\mathcal{S}=r+\dim\mathcal{S}^{\prime}<nd-2+\alpha(n-1,d)-(n-1)+1+\varepsilon=\alpha(n,d)-n+1+\varepsilon,$
@@
Hence $r^{\prime}=(n-1)d-1$ (remember that $\mathcal{S}^{\prime}$ is intransitive). Let us pick $\mathbb{F}$-linearly independent rows $L_{1}$ and $L_{2}$ in $\operatorname{\operatorname{M}}_{1,nd-r}(\mathbb{F})$. Set $\mathbf{R_{i}}:=L_{i}\mathbf{B_{1}}$ for $i\in\{1,2\}$. We have $\mathbf{R_{i}}\mathbf{C_{1}}=0$ with $\operatorname{\operatorname{rk}}\mathbf{C_{1}}=(n-1)d-1$ and the number of rows of $\mathbf{C_{1}}$ is at most (n1)d(n-1)d. Hence $\mathbf{R_{1}}$ and $\mathbf{R_{2}}$ are linearly dependent over $\mathbb{L}$. By Claim 5.3 (remembering that we assume r<nd1r<nd-1) we know that $\operatorname{\operatorname{sprk}}(\mathbf{R_{1}})>1$, and hence the Factorization Lemma yields a scalar $\alpha\in\mathbb{F}$ such that $\mathbf{R_{2}}=\alpha\mathbf{R_{1}}$. Thus $(L_{2}-\alpha L_{1})\mathbf{B_{1}}=0$ with $L_{2}-\alpha L_{1}\in\operatorname{\operatorname{M}}_{1,nd-r}(\mathbb{F})\smallsetminus\{0\}$, thereby contradicting Claim 5.2.
@@ 5.3 Deciphering the large spaces further
Now we proceed to prove the last point. So, now we assume that $\dim\mathcal{S}\geq\alpha(n,d)-n+2+\varepsilon$, and we prove that $\mathcal{S}$ has a nonzero alternator. Note that the assumptions of Claims 5.1 and 5.3 are satisfied. In particular $\mathcal{S}^{\prime}$ is target-reduced. Next,
$\dim\mathcal{S}^{\prime}=\dim\mathcal{S}-r\geq\alpha(n-1,d)-(n-1)+1+\varepsilon,$
and hence by induction we recover $\operatorname{\operatorname{trk}}(\mathcal{S}^{\prime})=(n-1)d-1$. It then follows from Proposition 4.4 that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S}^{\prime})\leq 1$.
@@
Since $\mathcal{S}^{\prime}$ is target-reduced, the identity $\mathbf{B_{1}}\mathbf{C_{1}}=0$ yields that $\mathbf{B_{1}}$ represents a nonzero alternator $b:U\times H\rightarrow\mathbb{F}$ of $\mathcal{S}^{\prime}$ (seen as a subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$). We deduce that $\operatorname{\operatorname{Alt}}(\mathcal{S}^{\prime})=\mathbb{F}b$, and hence the left radical
@@
does not depend on the choice of $b^{\prime}$ in $\operatorname{\operatorname{Alt}}(\mathcal{S})\smallsetminus\{0\}$ (in particular it is the left radical of $b$).
@@
Let $y\in U\smallsetminus\mathcal{L}$. The mappings $u\in\mathcal{S}$ such that $u(x)\in H$ do not all map $y$ into $\mathcal{S}x$.
@@
Assume on the contrary that $\operatorname{\operatorname{rk}}\mathbf{N}(y)<d$. Then there is a nonzero row $L=\begin{bmatrix}l_{1}\cdots l_{d}\end{bmatrix}\in\operatorname{\operatorname{M}}_{1,d}(\mathbb{F})$ such that $L\mathbf{N}(y)=0$. Denote by $\ell$ the $\mathbb{F}$-linear form on $V$ such that $\ell(e_{1})=\dots=\ell(e_{(n-1)d})=0$ and $\ell(e_{(n-1)d+i})=l_{i}$ for all $i\in\mathopen{[\![}1,d\mathclose{]\!]}$. This linear form is nonzero and its kernel includes $H$. The assumption $L\mathbf{N}(y)=0$ then shows that all the operators $u_{1},\dots,u_{(n-1)d}$ map $y$ into $\operatorname{\operatorname{Ker}}\ell$, and hence all the operators of $\mathcal{S}$ that map $x$ into $H$ also map $y$ into $\operatorname{\operatorname{Ker}}\ell$.
Next, we prove that $\operatorname{\operatorname{Ker}}\ell=(\mathcal{S}x)a^{-1}$ for some $a\in\mathbb{D}\smallsetminus\{0\}$. Choose indeed an $\mathbb{F}$-linear form $\ell^{\prime}$ on $V$ with kernel $\mathcal{S}x$. The linear forms $z\in V\mapsto\ell^{\prime}(za)$, with $a\in\mathbb{D}$, constitute a $d$-dimensional $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(V,\mathbb{F})$, and they all vanish on $H$. By orthogonality theory, it follows that they are exactly the $\mathbb{F}$-linear forms on $V$ whose kernel include $H$. In particular there exists $a\in\mathbb{D}$ such that $\forall z\in V,\;\ell(z)=\ell^{\prime}(za)$, and obviously $a\neq 0$. It follows that
$\operatorname{\operatorname{Ker}}\ell=(\operatorname{\operatorname{Ker}}\ell^{\prime})a^{-1}=(\mathcal{S}x)a^{-1}=\mathcal{S}(xa^{-1}).$
Now the trick is to use the fact that $xa^{-1}$ satisfies exactly the same assumptions as the vector $x$ we started from. Indeed $\operatorname{\operatorname{rk}}(\widehat{xa^{-1}})=\dim_{\mathbb{F}}\bigl{(}\mathcal{S}(xa^{-1})\bigr{)}=nd-1$, the operators in $\mathcal{S}$ that vanish at $x$ are also the ones that vanish at $xa^{-1}$, and $H$ is the $\mathbb{D}$-linear hyperplane of $V$ that is included in $\mathcal{S}(xa^{-1})$. Hence Claim 5.5 applies when $x$ is replaced with $xa^{-1}$. Every $u\in\mathcal{S}$ that maps $x$ into $H$ also maps $xa^{-1}$ into $H$ because $u$ is $\mathbb{D}$-linear, and hence there is at least one such operator that does not map $y$ into $\mathcal{S}(xa^{-1})$. This contradicts the first step of the present proof. We conclude that $\operatorname{\operatorname{rk}}\mathbf{N}(y)\geq d$, which completes the proof. ∎
5.4 Constructing a nonzero alternator of $\mathcal{S}$
Now everything is in place for us to produce a nonzero alternator of $\mathcal{S}$.
@@ 5.4 Constructing a nonzero alternator of $\mathcal{S}$
Now, we consider the new row $\mathbf{R}:=\mathbf{B}(\mathbf{A}-\mathbf{a}I_{nd-1})$, which can be written $\begin{bmatrix}[0]_{1\times(n-1)d}&\mathbf{R_{2}}\end{bmatrix}$. We also write
@@
Since it was known from the start that $\mathcal{S}$ is target-reduced, the process explained in the end of Section 4 yields a nonzero alternator $\widetilde{b}:U\times V\rightarrow\mathbb{F}$ of $\mathcal{S}$, associated with the nonzero catcher $\begin{bmatrix}\mathbf{B}&-\mathbf{a}\end{bmatrix}$.
Finally, since $\mathcal{S}$ has transitive rank nd1nd-1 and is target-reduced, we know from Proposition 4.4 that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})\leq 1$, and we conclude that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$.
@@ 5.5 Operator spaces with $1$-dimensional alternator spaces
We will prove the following general result, which avoids any dimensional assumption on $\mathcal{S}$ and focuses on the condition regarding the structure of the alternator space.
@@
Let $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$, and let $b\in\operatorname{\operatorname{Alt}}(\mathcal{S})\smallsetminus\{0\}$. Then:
@@
Then $a\mapsto[c\mapsto c^{a}]$ defines an $\mathbb{F}$-quadratic mapping $P:\mathbb{D}\rightarrow\operatorname{\operatorname{End}}_{\mathbb{F}}(\mathcal{B}(U\times V,\mathbb{F}))$, with $P(aa^{\prime})=P(a)\circ P(a^{\prime})$ for all $(a,a^{\prime})\in\mathbb{D}^{2}$, i.e., $P$ is multiplicative. Moreover, $P(a)$ leaves $\operatorname{\operatorname{Alt}}(\mathcal{S})$ invariant for all $a\in\mathbb{D}$, as for all $c\in\operatorname{\operatorname{Alt}}(\mathcal{S})$ we observe that
$\forall u\in\mathcal{S},\;\forall x\in U,\;c(xa,u(x)a)=c(xa,u(xa))=0.$
Since $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$, we deduce that $b$ is an eigenvector of $P(a)$ for all $a\in\mathbb{D}$, and by denoting by $q(a)\in\mathbb{F}$ the corresponding eigenvalue we obtain an $\mathbb{F}$-quadratic form $q:\mathbb{D}\rightarrow\mathbb{F}$. Obviously $\operatorname{\operatorname{rk}}b^{a}=\operatorname{\operatorname{rk}}b$ for all $a\in\mathbb{D}^{\times}$ (because $\mathbb{D}$ is a division ring), and in particular $q$ is nonisotropic. Finally, since $P$ is multiplicative we obtain that $q$ is multiplicative.
Therefore, we are in the position to use Theorem 2.3: the algebra $\mathbb{D}$ has quadratic type over $\mathbb{F}$. Let us now choose a pair $(\sigma,e)$ that is associated with $(\mathbb{D},\mathbb{F})$. Then we also know that $q(a)=a\sigma(a)$ for all $a\in\mathbb{D}$. Moreover $e$ satisfies the identity $\forall(x,y)\in\mathbb{D}^{2},\;e(yx)=e(xy)$, which will ease things up from now on. As seen in Section 2.1, since $e$ is nonzero we recover a unique $\mathbb{F}$-bilinear and right-$\mathbb{D}$-linear form $B$ on $U\times V$ such that
@@
Now we can complete the proof of point (c) of Theorem 1.20. Let $\mathcal{S}$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ with $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$ and $|\mathbb{F}|\geq nd$. By Proposition 5.1 we have $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})=1$.
Choose $b\in\operatorname{\operatorname{Alt}}(\mathcal{S})\smallsetminus\{0\}$. We need to prove that $b$ is right-nondegenerate. Set $R:=\operatorname{\operatorname{Rrad}}(b)$, which by point (c) of Proposition 5.2 is a $\mathbb{D}$-linear subspace of $V$. Set $t:=\dim_{\mathbb{D}}R\in\mathopen{[\![}0,n-1\mathclose{]\!]}$, and denote by $\pi:V\twoheadrightarrow V/R$ the standard projection, which is $\mathbb{D}$-linear. The bilinear form $b$ induces a right-nondegenerate $\mathbb{F}$-bilinear form $\overline{b}$ on $U\times(V/R)$. The space $\pi\mathcal{S}$ is an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/R)$, and it is obvious that $\overline{b}$ is an alternator of it. Since $\overline{b}$ is right-nondegenerate, we know from Proposition 1.18 that $\pi\mathcal{S}$ is deeply intransitive. Besides, the kernel of $u\in\mathcal{S}\mapsto\pi\circ u$ is precisely $\mathcal{S}^{R}$, which is deeply intransitive. Hence, by combining the rank theorem with point (a) of Theorem 1.20, we find
$\dim(\mathcal{S})=\dim(\pi\mathcal{S})+\dim(\mathcal{S}^{R})\leq\alpha(n-t,d)+\alpha(t,d)=\alpha(n,d)-dt(n-t).$
It $t>0$ then $dt(n-t)\geq d(n-1)$, and we observe that $d(n-1)>n-\max(4-d,2)$ because obviously $(d-1)(n-1)\geq 0>1-\max(4-d,2)$. This runs counter to our starting assumption on the dimension of $\mathcal{S}$.

### v2

@@ 5.2 The transitive rank
Denote by $r^{\prime}$ the rank of $\mathbf{C_{1}}$: it is also the transitive rank of $S^{\prime}$. If $r^{\prime}<(n-1)d-1$, then by induction $\dim S^{\prime}<\alpha(n-1,d)-(n-1)+1+\varepsilon$ and hence
$\dim S=r+\dim S^{\prime}<nd-2+\alpha(n-1,d)-(n-1)+1+\varepsilon=\alpha(n,d)-n+1+\varepsilon,$
@@
Hence $r^{\prime}=(n-1)d-1$ (remember that $S^{\prime}$ is intransitive). Let us pick $\mathbb{F}$-linearly independent rows $L_{1}$ and $L_{2}$ in $\operatorname{\operatorname{M}}_{1,nd-r}(\mathbb{F})$. Set $\mathbf{R_{i}}:=L_{i}\mathbf{B_{1}}$ for $i\in\{1,2\}$. We have $\mathbf{R_{i}}\mathbf{C_{1}}=0$ with $\operatorname{\operatorname{rk}}\mathbf{C_{1}}=(n-1)d-1$ and the number of rows of $\mathbf{C_{1}}$ is at most (n1)d(n-1)d. Hence $\mathbf{R_{1}}$ and $\mathbf{R_{2}}$ are linearly dependent over $\mathbb{L}$. By Claim 5.3 (remembering that we assume r<nd1r<nd-1) we know that $\operatorname{\operatorname{sprk}}(\mathbf{R_{1}})>1$, and hence the Factorization Lemma yields a scalar $\alpha\in\mathbb{F}$ such that $\mathbf{R_{2}}=\alpha\mathbf{R_{1}}$. Thus $(L_{2}-\alpha L_{1})\mathbf{B_{1}}=0$ with $L_{2}-\alpha L_{1}\in\operatorname{\operatorname{M}}_{1,nd-r}(\mathbb{F})\smallsetminus\{0\}$, thereby contradicting Claim 5.2.
@@ 5.3 Deciphering the large spaces further
Now we proceed to prove the last point. So, now we assume that $\dim S\geq\alpha(n,d)-n+2+\varepsilon$, and we prove that $S$ has a nonzero alternator. Note that the assumptions of Claims 5.1 and 5.3 are satisfied. In particular $S^{\prime}$ is target-reduced. Next,
$\dim S^{\prime}=\dim S-r\geq\alpha(n-1,d)-(n-1)+1+\varepsilon,$
and hence by induction we recover $\operatorname{\operatorname{trk}}(S^{\prime})=(n-1)d-1$. It then follows from Proposition 4.4 that $\dim\operatorname{\operatorname{Alt}}(S^{\prime})\leq 1$.
@@
Since $S^{\prime}$ is target-reduced, the identity $\mathbf{B_{1}}\mathbf{C_{1}}=0$ yields that $\mathbf{B_{1}}$ represents a nonzero alternator $b:U\times H\rightarrow\mathbb{F}$ of $S^{\prime}$ (seen as a subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,H)$). We deduce that $\operatorname{\operatorname{Alt}}(S^{\prime})=\mathbb{F}b$, and hence the left radical
@@
does not depend on the choice of $b^{\prime}$ in $\operatorname{\operatorname{Alt}}(S)\smallsetminus\{0\}$ (in particular it is the left radical of $b$).
@@
Let $y\in U\smallsetminus\mathcal{L}$. The mappings $u\in S$ such that $u(x)\in H$ do not all map $y$ into $Sx$.
@@
Assume on the contrary that $\operatorname{\operatorname{rk}}\mathbf{N}(y)<d$. Then there is a nonzero row $L=\begin{bmatrix}l_{1}\cdots l_{d}\end{bmatrix}\in\operatorname{\operatorname{M}}_{1,d}(\mathbb{F})$ such that $L\mathbf{N}(y)=0$. Denote by $\ell$ the $\mathbb{F}$-linear form on $V$ such that $\ell(e_{1})=\dots=\ell(e_{(n-1)d})=0$ and $\ell(e_{(n-1)d+i})=l_{i}$ for all $i\in\mathopen{[\![}1,d\mathclose{]\!]}$. This linear form is nonzero and its kernel includes $H$. The assumption $L\mathbf{N}(y)=0$ then shows that all the operators $u_{1},\dots,u_{(n-1)d}$ map $y$ into $\operatorname{\operatorname{Ker}}\ell$, and hence all the operators of $S$ that map $x$ into $H$ also map $y$ into $\operatorname{\operatorname{Ker}}\ell$.
Next, we prove that $\operatorname{\operatorname{Ker}}\ell=(Sx)a^{-1}$ for some $a\in\mathbb{D}\smallsetminus\{0\}$. Choose indeed an $\mathbb{F}$-linear form $\ell^{\prime}$ on $V$ with kernel $Sx$. The linear forms $z\in V\mapsto\ell^{\prime}(za)$, with $a\in\mathbb{D}$, constitute a $d$-dimensional $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{F}}(V,\mathbb{F})$, and they all vanish on $H$. By orthogonality theory, it follows that they are exactly the $\mathbb{F}$-linear forms on $V$ whose kernel include $H$. In particular there exists $a\in\mathbb{D}$ such that $\forall z\in V,\;\ell(z)=\ell^{\prime}(za)$, and obviously $a\neq 0$. It follows that
$\operatorname{\operatorname{Ker}}\ell=(\operatorname{\operatorname{Ker}}\ell^{\prime})a^{-1}=(Sx)a^{-1}=S(xa^{-1}).$
Now the trick is to use the fact that $xa^{-1}$ satisfies exactly the same assumptions as the vector $x$ we started from. Indeed $\operatorname{\operatorname{rk}}(\widehat{xa^{-1}})=\dim_{\mathbb{F}}\bigl(S(xa^{-1})\bigr)=nd-1$, the operators in $S$ that vanish at $x$ are also the ones that vanish at $xa^{-1}$, and $H$ is the $\mathbb{D}$-linear hyperplane of $V$ that is included in $S(xa^{-1})$. Hence Claim 5.5 applies when $x$ is replaced with $xa^{-1}$. Every $u\in S$ that maps $x$ into $H$ also maps $xa^{-1}$ into $H$ because $u$ is $\mathbb{D}$-linear, and hence there is at least one such operator that does not map $y$ into $S(xa^{-1})$. This contradicts the first step of the present proof. We conclude that $\operatorname{\operatorname{rk}}\mathbf{N}(y)\geq d$, which completes the proof. ∎
5.4 Constructing a nonzero alternator of $S$
Now everything is in place for us to produce a nonzero alternator of $S$.
@@ 5.4 Constructing a nonzero alternator of $S$
Now, we consider the new row $\mathbf{R}:=\mathbf{B}(\mathbf{A}-\mathbf{a}I_{nd-1})=\begin{bmatrix}[0]_{1\times(n-1)d}&\mathbf{R_{2}}\end{bmatrix}$. We also write
@@
Since it was known from the start that $S$ is target-reduced, the process explained in the end of Section 4 yields a nonzero alternator $\widetilde{b}:U\times V\rightarrow\mathbb{F}$ of $S$, associated with the nonzero catcher $\begin{bmatrix}\mathbf{B}&-\mathbf{a}\end{bmatrix}$.
Finally, since $S$ has transitive rank nd1nd-1 and is target-reduced, we know from Proposition 4.4 that $\dim\operatorname{\operatorname{Alt}}(S)\leq 1$, and we conclude that $\dim\operatorname{\operatorname{Alt}}(S)=1$.
@@ 5.5 Operator spaces with $1$-dimensional alternator spaces
We will prove the following general result, which avoids any dimensional assumption on $S$ and focuses on the condition regarding the structure of the alternator space.
@@
Let $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\operatorname{\operatorname{Alt}}(S)$ has dimension $1$, and let $b\in\operatorname{\operatorname{Alt}}(S)\smallsetminus\{0\}$. Then:
@@
Then $a\mapsto[c\mapsto c^{a}]$ defines an $\mathbb{F}$-quadratic mapping $P:\mathbb{D}\rightarrow\operatorname{\operatorname{End}}_{\mathbb{F}}(\mathcal{B}(U\times V,\mathbb{F}))$, with $P(aa^{\prime})=P(a)\circ P(a^{\prime})$ for all $(a,a^{\prime})\in\mathbb{D}^{2}$, i.e., $P$ is multiplicative. Moreover, $P(a)$ leaves $\operatorname{\operatorname{Alt}}(S)$ invariant for all $a\in\mathbb{D}$, as for all $c\in\operatorname{\operatorname{Alt}}(S)$ we observe that
$\forall u\in S,\;\forall x\in U,\;c(xa,u(x)a)=c(xa,u(xa))=0.$
Since $\operatorname{\operatorname{Alt}}(S)$ has dimension $1$, we deduce that $b$ is an eigenvector of $P(a)$ for all $a\in\mathbb{D}$, and by denoting by $q(a)\in\mathbb{F}$ the corresponding eigenvalue we obtain an $\mathbb{F}$-quadratic form $q:\mathbb{D}\rightarrow\mathbb{F}$. Obviously $\operatorname{\operatorname{rk}}b^{a}=\operatorname{\operatorname{rk}}b$ for all $a\in\mathbb{D}^{\times}$ (because $\mathbb{D}$ is a division ring), and in particular $q$ is nonisotropic. Finally, since $P$ is multiplicative we obtain that $q$ is multiplicative.
Therefore, we are in the position to use Theorem 2.3: the algebra $\mathbb{D}$ has quadratic type over $\mathbb{F}$. Let us now choose a pair $(\sigma,e)$ that is associated with $(\mathbb{D},\mathbb{F})$. Then we also know that $q(a)=a\sigma(a)$ for all $a\in\mathbb{D}$. Moreover $e$ satisfies the identity $\forall(x,y)\in\mathbb{D}^{2},\;e(yx)=e(xy)$, which will ease things up from now on. As seen in Section 2, since $e$ is nonzero we recover a unique $\mathbb{F}$-bilinear and right-$\mathbb{D}$-linear form $B$ on $U\times V$ such that
@@
Now we can complete the proof of point (c) of Theorem 1.22. Let $S$ be a deeply intransitive $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ with $\dim S\geq\alpha(n,d)-n+\max(4-d,2)$ and $|\mathbb{F}|\geq nd$. By Proposition 5.1 we have $\dim\operatorname{\operatorname{Alt}}(S)=1$.
Choose $b\in\operatorname{\operatorname{Alt}}(S)\smallsetminus\{0\}$. We need to prove that $b$ is right-nondegenerate. Set $R:=\operatorname{\operatorname{Rrad}}(b)$, which by point (c) of Proposition 5.2 is a $\mathbb{D}$-linear subspace of $V$. Set $t:=\dim_{\mathbb{D}}R\in\mathopen{[\![}0,n-1\mathclose{]\!]}$, and denote by $\pi:V\twoheadrightarrow V/R$ the standard projection, which is $\mathbb{D}$-linear. The bilinear form $b$ induces a right-nondegenerate $\mathbb{F}$-bilinear form $\overline{b}$ on $U\times(V/R)$. The space $\pi S$ is an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V/R)$, and it is obvious that $\overline{b}$ is an alternator of it. Since $\overline{b}$ is right-nondegenerate, we know from Proposition 1.20 that $\pi S$ is deeply intransitive. Besides, the kernel of $u\in S\mapsto\pi\circ u$ is precisely $S^{R}$, which is deeply intransitive. Hence, by combining the rank theorem with point (a) of Theorem 1.22, we find
$\dim(S)=\dim(\pi S)+\dim(S^{R})\leq\alpha(n-t,d)+\alpha(t,d)=\alpha(n,d)-dt(n-t).$
It $t>0$ then $dt(n-t)\geq d(n-1)$, and we observe that $d(n-1)>n-\max(4-d,2)$ because obviously $(d-1)(n-1)\geq 0>1-\max(4-d,2)$. This runs counter to our starting assumption on the dimension of $S$.